# Batched Self-Consistency Improves LLM Relevance Assessment and Ranking

### v1

@@ Abstract
Given some information need, Large Language Models (LLMs) are increasingly used for candidate text relevance assessment, typically using a one-by-one pointwise (PW) strategy where each LLM call evaluates one candidate at a time. Meanwhile, it has been shown that LLM performance can be improved through self-consistency: prompting the LLM to do the same task multiple times (possibly in perturbed ways) and then aggregating the responses. To take advantage of self-consistency, we hypothesize that batched PW strategies, where multiple passages are judged in one LLM call, are better suited than one-by-one PW methods since a larger input context can induce more diverse LLM sampling across self-consistency calls. We first propose several candidate batching strategies to create prompt diversity across self-consistency calls through subset reselection and permutation. We then test our batched PW methods on relevance assessment and ranking tasks against one-by-one PW and listwise LLM ranking baselines with and without self-consistency, using three passage retrieval datasets and GPT-4o, Claude Sonnet 3, and Amazon Nova Pro. We find that batched PW methods outperform all baselines, and show that batching can greatly amplify the positive effects of self-consistency. For instance, on our legal search dataset, GPT-4o one-by-one PW ranking NDCG@10 improves only from 44.9% to 46.8% without self-consistency vs. with 15 self-consistency calls, while batched PW ranking improves from 43.8% to 51.3%.
@@ 1 Introduction
LLM relevance assessment and ranking have become two foundational tasks for agentic artificial intelligence (AI) and retrieval-augmented generation (RAG) systems Wang et al. (2024). Given an information need, existing LLM candidate relevance assessment methods typically use a one-by-one pointwise (PW) strategy where each LLM call evaluates one candidate at a time Thomas et al. (2024); Upadhyay et al. (2024). While such methods can avoid positional biases due to candidate order, they prevent information sharing between multiple candidates and can be computationally expensive, requiring a separate LLM call for each candidate. For LLM ranking, several techniques exist including PW, listwise (LW) Tang et al. (2024); Ma et al. (2023); Sun et al. (2023), and pairwise Qin et al. (2024); Liusie et al. (2024) approaches, but only PW methods can rank and generate absolute relevance judgments.
Meanwhile, a simple technique for improving LLM output generation quality has emerged called self-consistency Wang et al. (2023) which prompts the LLM to do the same task multiple times (possibly with prompt perturbations) and then aggregates the outputs – leveraging the variability in LLM decoding. Motivated by these observations, we propose that batched PW scoring, where multiple passages receive relevance scores in one LLM call, is especially well suited for self-consistency methods. Compared to one-by-one PW appraoches, we hypothesize that the larger context provided by batched PW methods can induce more helpful variation between LLM self-consistency calls, as well as benefit from diverse prompts created with various candidate subsets and permutations.
We evaluate these batched PW methods against one-by-one PW and LW techniques with and without self-consistency using GPT-4o, Claude Sonnet 3, and Amazon Nova Pro on passage relevance assessment and ranking tasks from three datsets. Our contributions include:
  • We propose several batching strategies (Figure 1) to diversify LLM self-consistency prompts using various candidate subsets and permutations.
  • We show that batching can greatly amplify the positive effects of self-consistency. For instance, on legal search, GPT-4o one-by-one PW ranking NDCG@10 improves from 44.9% without self-consistency to 46.8% with self-consistency – while the best batched PW ranking method improves from 43.8% to 51.3%, respectively.
  • We find that large batches can introduce harmful position biases in all three LLMs, but these can be successfully mitigated by using smaller batches, shuffling, and self-consistency.
  • We observe that batched self-consistency PW methods outperform all LW and one-by-one PW methods with self-consistency.
@@ 2.1 Text Relevance Assessment and Ranking
Given a query $q$, both non-generative and generative approaches are widely used to rank or score candidate text spans in some collection $\{p_{1},\cdots,p_{D}\}$ based on relevance to $q$. Non-neural, sparse methods such as TF-IDF Salton et al. (1975) and its probabilistic variant BM25 Robertson et al. (2009) rely on syntactic token matches, limiting their ability to capture semantic similarity. Encoder-only LLM methods broadly include bi-encoders Izacard et al. (2021); Gao and Callan (2021) which score using a similarity function (e.g. cosine similarity) between separately embedded queries and passages, and cross-encoders Nogueira and Cho (2019); Zhuang et al. (2023) which jointly embed queries and passages to predict a relevance score. While these foundational methods are critical for retrieving initial candidate lists from large corpora, none of them are able to benefit from self-consistency since they do not use stochastic decoding or prompting.
Generative LLM-based methods typically consist of PW, pairwise, and LW strategies. Standard one-by-one PW and LW techniques (with and without self consistency) are baselines discussed further in sections 3.1.1 and 3.1.2, respectively. Pairwise rankers Qin et al. (2024); Liu et al. (2024) ask an LLM which passage out of a pair $(p_{i},p_{j})$ is more relevant, but need a quadratic number of LLM calls relative to the number of candidates. There are also bubble-sort sliding-window based LW LLM rankers Ma et al. (2023); Sun et al. (2023) requiring sequential LLM calls, but we instead focus on fully parallelizable methods.
@@ 2.2 LLM Self-Consistency
LLM self-consistency Wang et al. (2023) involves prompting an LLM to do the same task multiple times (potentially with perturbed prompts) and then aggregating the outputs, improving response quality by leveraging the stochasticity in LLM decoding. Multiple variations exist Li et al. (2024), including using perturbations of few-shot examples to diversify prompts Lu et al. (2022), and using heterogeneous LLM ensembles Wan et al. (2024). Recent work Tang et al. (2024); Hou et al. (2024) reported that self-consistency can improve LW LLM ranking, as discussed further in Section 3.1.2
@@ 3 Methodology
In all PW methods, each LLM call generates a relevance score $s_{q,p}\in\mathbb{R}$ between $q$ and each passage $p\in L^{q}$. For ranking, passages are sorted in descending score order with ties broken using the order of the initial list $L^{q}$. We use the following 0-3 scale from the UMBRELLA open-source Bing prompt Upadhyay et al. (2024), with our full prompt shown in Appendix C:
@@
All our self-consistency methods include each passage in exactly $m$ LLM calls. In our PW self-consistency methods, each passage thus receives $m$ scores $\{s^{1}_{q,p},\cdots,s^{m}_{q,p}\}$, which are aggregated into a final score $s_{q,p}$ by taking the mean. [^1] For ranking with integer 0-3 scores, mean aggregation reduces the number of ties compared to majority voting. Similarly, our LW methods aggregate $m$ output lists by minimizing Kendall-Tau distance, as described further in Sec. 3.1.2.
Figure 2: Effect of increasing self-consistency calls/passage ($m$) on PW relevance assessment quality (Legal Search, Shallow). For all LLMs at $m=1$ (no self-consistency), one-by-one PW is competitive, but by $m=15$ it underperforms batched PW by 5-10% AUC-PR. This shows that batching amplifies the benefits of self-consistency, likely due to broader input contexts.
Figure 3: Legal Search, Deep (90 Total Psgs): effect of increasing $m$ on relevance assessment quality. The sub-batched methods (30 psgs/batch) perform very well, with the shuffling variants (STB and BTS) doing best. The large batch (90 psgs/batch) all-in-one methods perform poorly for this range of $m$, addressed further in RQ2.
Table 1: AUCm, representing the AUC-PR at $m$ self-consistency calls/passage, for PW relevance assessment methods at $m=1$ (no self-consistency) and $m=15$. Increasing $m$ improves all PW methods, but the batched PW methods improve more, becoming the best methods at $m=15$, likely due to their larger prompt contexts. The highest AUC-PR for each $m$ and LLM is in bold.
Legal Search DL-19 Covid
GPT-4o Sonnet Nova GPT-4o Sonnet Nova GPT-4o Sonnet Nova
Psg/ Psg AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/ AUC1/
Batch Order AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15 AUC15
  Shallow (30 psg) 1 32 / 41 24 / 28 24 / 29 39 / 52 27 / 31 29 / 33 70 / 78 68 / 73 68 / 77
12
10 Init. 31 / 44 22 / 27 27 / 33 34 / 51 25 / 33 32 / 40 70 / 77 63 / 70 68 / 74
(sub-batch) STB 31 / 46 22 / 38 25 / 39 39 / 57 25 / 41 30 / 47 69 / 80 65 / 77 68 / 79
BTS 31 / 45 23 / 35 25 / 36 39 / 55 25 / 41 32 / 46 69 / 79 63 / 75 65 / 78
12
30 Init. 32 / 45 23 / 27 27 / 37 46 / 55 26 / 35 30 / 44 69 / 78 63 / 69 63 / 74
(all psgs) Shuf. 30 / 45 20 / 32 24 / 38 37 / 58 24 / 41 31 / 53 69 / 80 58 / 74 65 / 77
12
  Deep (90 psg) 1 20 / 29 15 / 20 15 / 20 34 / 51 23 / 28 25 / 31 63 / 72 62 / 68 62 / 70
12
30 Init. 20 / 34 13 / 21 18 / 28 30 / 49 20 / 30 29 / 40 64 / 75 55 / 67 60 / 72
(sub-batch) STB 20 / 36 14 / 27 17 / 31 32 / 52 19 / 40 23 / 42 64 / 77 56 / 73 60 / 75
BTS 22 / 36 14 / 26 17 / 29 29 / 52 18 / 37 25 / 44 64 / 77 55 / 72 60 / 75
12
90 Init. 16 / 26 11 / 16 13 / 20 22 / 48 16 / 33 23 / 37 52 / 67 48 / 57 45 / 54
(all psg) Shuf. 14 / 29 11 / 21 13 / 20 22 / 50 15 / 42 19 / 44 53 / 73 45 / 65 46 / 60
12
@@ 3.1.1 One-by-one Pointwise Methods
The first pane of Figure 1 shows one-by-one PW methods Thomas et al. (2024); Upadhyay et al. (2024); Törnberg (2024), where each passage is scored against the query in a separate LLM call. While evaluating one candidate at a time avoids inducing candidate order position biases, it also prevents the LLM from seeing potentially helpful context in other passages. Further, each passage score requires a separate LLM call, which typically leads to a very large number of calls – especially with self-consistency.
@@ 3.1.2 Listwise Ranking Methods
LW ranking methods Ma et al. (2023); Tang et al. (2024); Hou et al. (2024) are shown in the second pane of Figure 1.
@@
Standard LW ranking instructs an LLM to rerank an input passage list in order of relevance with respect to $q$, with our LW prompt shown in Appendix C. While the LLM sees all available passage context during a single inference, no absolute relevance judgments are produced.
@@
LW ranking with self-consistency Tang et al. (2024); Hou et al. (2024) involves $m$ reranking calls followed by a rank aggregation of $m$ output passage lists, which can be done by minimizing the Kendall-Tau distance with the lists. Multiple exact and approximate aggregation techniques exist, with our experiments using the exact Kemeny rank aggregation linear program (LP) of Tang et al. [^2] https://github.com/castorini/perm-sc We test two LW self-consistency variants:
@@ 3.2 Batched PW Methods
The right half of Figure 1 shows batched PW methods in which multiple passages are jointly scored in each LLM call. This allows an LLM to see context from multiple passages at the same time and reduces the total number of LLM calls required compared to one-by-one methods. We study the effects of diversifying passage subsets and permutations through several batching strategies, including all-in-one batching (c.f. Sec. 3.2.1) and sub-batching (c.f. Sec. 3.2.2).
@@ 3.2.1 All-in-one
All-in-one PW methods prompt the LLM to score all passages in a single batch, [^3] The LLM context window must be large enough to fit all candidate passages, otherwise sub-batching (c.f. Sec 3.2.2) is required. maximizing the context available to the model but also presenting it with the most complex task. We test two passage ordering strategies:
@@
    Shuffled: The passage list is fully shuffled before each self-consistency LLM call, giving $m$ calls with $m$ random passage perturbations.
@@ 3.2.2 Sub-batched
    Initial Order: To create $B$ batches, the initial list is partitioned into $B$ non-overlapping intervals while maintaining its order (e.g., $[p_{1},\cdots,p_{30}]$rightarrow $[p_{1}\cdots,p_{10}],$ $[p_{11}\cdots,p_{20}],$ $[p_{21},\cdots,p_{30}]$, where $p_{i}$ is the $i$’th passage in the initial list). As shown in the top right of Figure 1, self-consistency calls for each sub-batch are identical, giving this sub-batching strategy the least diversity in LLM sampling.
@@
    Shuffled-then-Batched (STB): The passage list is fully shuffled and then split into $B$ batches before each LLM call. When self-consistency is used, STB creates the most diverse range of passage subsets and permutations across LLM calls.
@@
    Batched-then-Shuffled (BTS): The initial list is first divided into $B$ intervals, as in Initial Order Sub-batching above, and then each batch $b$ is fully shuffled before each LLM call. While BTS attempts to mitigate position bias through shuflling, the passage mixture for a given batch remains constant across self-consistency calls, making its LLM sampling less diverse than STB.

### v2

@@ Abstract
LLM query-passage relevance assessment is typically studied using a one-by-one pointwise (PW) strategy where each LLM call judges one passage at a time. However, this strategy requires as many LLM calls as there are passages while also preventing information sharing between passages. We thus hypothesize that batched PW methods, which evaluate multiple passages per LLM call, can improve not only efficiency but also judgment quality – by enabling content from multiple passages to be seen jointly. Moreover, batched PW methods may be better suited to harness the test-time scaling benefits of self-consistency -— the ensembling technique of repeating (potentially perturbed) LLM tasks in parallel and aggregating results -— since batching can naturally enable prompt diversification through varied batch permutations and compositions to create more robust ensembles. We evaluate several batched PW methods against one-by-one PW and listwise ranking baselines on LLM relevance assessment and ranking tasks, using three passage retrieval datasets and GPT-4o, Claude Sonnet 3, and Amazon Nova Pro. We show that batching can greatly amplify self-consistency benefits, making batched PW methods achieve the best performance while often reducing latency by an order of magnitude or more compared to one-by-one PW methods. For instance, on legal search, batched PW ranking with GPT-4o improves from 43.8% to 51.3% NDCG@10 when using 1 vs. 15 self-consistency calls, compared to one-by-one PW ranking improving from 44.9% to 46.8% and being 15.3x slower.
@@ 1 Introduction
Given a query and a set of candidate passages, LLM query-passage relevance assessment has become a foundational task for agentic artificial intelligence (AI) and retrieval-augmented generation (RAG) Wang et al. (2024). Existing work has focused on a one-by-one pointwise (PW) scoring strategy where each LLM call judges one passage at a time against the query Sachan et al. (2022); Zhuang et al. (2024); Thomas et al. (2024); Upadhyay et al. (2024); Törnberg (2024); Ma et al. (2024). However, this strategy does not allow content from multiple candidate passages to be seen jointly, and can be computationally expensive as it requires as many LLM calls as there are passages. In this work, we hypothesize that batched PW scoring – evaluating multiple candidates in a single LLM call – can improve both judgment quality and computational efficiency, since content from multiple passages can be seen jointly while requiring fewer LLM calls.
Further, we ask how LLM relevance assessment methods can leverage the test-time scaling benefits of self-consistency Wang et al. (2023) – a parallelizable ensembling technique where the same LLM task is repeated, potentially with perturbations, and the results are aggregated. To collect multiple scores per passage, one-by-one methods are generally limited to repeating identical LLM calls, relying only on LLM stochasticity to produce different scores. Batched PW methods, however, can naturally perturb the contexts in which scores for a given candidate are generated by using different co-candidate subsets and permutations across batches, as shown in Figure 1. Moreover, autoregressively generating a sequence of scores, as opposed to only one score at a time, could diversify scoring further. We thus conjecture that batching may create more robust self-consistency ensembles that exhibit stronger test time scaling due to more diversification across ensemble components.
We conduct experiments on the tasks of passage relevance assessment (i.e., predicting the relevance label of a passage) and ranking, using three passage retrieval datasets and three LLMs (GPT-4o, Clause Sonnet 3, and Amazon Nova Pro). We compare batched PW methods against one-by-one PW and listwise (LW) Ma et al. (2023); Tang et al. (2024); Hou et al. (2024) baselines – for each method, we also investigate the effects of self-consistency for ensemble sizes from 1-15 components. Our contributions include:
  • We show that batched PW methods can improve not only efficiency but also judgment quality by enabling content from multiple passages to be seen jointly.
  • We propose several batched PW strategies to diversify self-consistency ensembles via various candidate subsets and permutations.
  • We find that batching produces stronger self-consistency ensembles with faster test-time scaling. For instance, on legal search, batched PW ranking with GPT-4o improves from 43.8% to 51.3% NDCG@10 when using 1 vs. 15 self-consistency calls, compared to one-by-one PW ranking which improves from 44.9% to 46.8%.
  • We show that batching can reduce latency by an order of magnitude or more when the number of available LLMs does not exceed the number of ensemble components.
@@ 2.1 Text Relevance Assessment and Ranking
Given a query $q$, both non-generative and generative approaches are widely used to rank or score candidate text spans in some collection $\{p_{1},\cdots,p_{D}\}$ based on relevance to $q$. Non-neural, sparse methods such as TF-IDF Salton et al. (1975) and its probabilistic variant BM25 Robertson et al. (2009) rely on syntactic token matches, limiting their ability to capture semantic similarity. Encoder-only LLM methods broadly include bi-encoders Izacard et al. (2021); Gao and Callan (2021) which score using a similarity function (e.g. cosine similarity) between separately embedded queries and passages, and cross-encoders Nogueira and Cho (2019); Zhuang et al. (2023) which jointly embed queries and passages to predict a relevance score. While these foundational methods are critical for retrieving initial candidate lists from large corpora, none of them are able to benefit from self-consistency since they do not use prompting or stochastic decoding.
Generative LLM-based methods typically consist of PW, pairwise, and LW strategies. Standard one-by-one PW and LW techniques are baselines discussed further in sections 3.1.1 and 3.1.2, respectively. Pairwise rankers Qin et al. (2024); Liu et al. (2024) ask an LLM which passage out of a pair $(p_{i},p_{j})$ is more relevant, but need a quadratic number of LLM calls relative to the number of candidates. There are also bubble-sort sliding-window LW LLM rankers Ma et al. (2023); Sun et al. (2023); Pradeep et al. (2023a, b) which achieve beneficial test-time scaling effects by using multiple sequential LLM calls to iteratively rank overlapping intervals of the initial list. We leave an exploration of sequential test-time scaling for batched PW methods for future work, focusing instead on fully parallelizable self-consistency ensembles.
@@ 2.2 LLM Self-Consistency
LLM self-consistency Wang et al. (2023) is a parallelizable ensembling technique that involves performing the same LLM task multiple times and aggregating the outputs. The variation between outputs can occur not only because of stochastic response generation but also from perturbations to the prompt. Many approaches exist Zhang et al. (2024), including diversifying few-shot examples across self-consistency calls Lu et al. (2022), using token entropy to weigh or prune components Kang et al. (2025); Fu et al. (2025), creating ensembles of heterogenous LLMs Wan et al. (2024), and recent work on LW ranking that uses different passage perturbations across calls Tang et al. (2024); Hou et al. (2024), as discussed further in Section 3.1.2.
@@ 3 Methodology
In all PW methods, each LLM call generates a relevance score $s_{q,p}\in\mathbb{R}$ between $q$ and each passage $p\in L^{q}$. For ranking, passages are sorted in descending score order with ties broken using the order of the initial list $L^{q}$. We use the following 0-3 scale from the UMBRELLA open-source Bing prompt Upadhyay et al. (2024):
@@
All our self-consistency methods include each passage in exactly $m$ LLM calls. In our PW self-consistency methods, each passage thus receives $m$ scores $\{s^{1}_{q,p},\cdots,s^{m}_{q,p}\}$, which are aggregated into a final score $s_{q,p}$ by taking the mean. [^1] For ranking with integer 0-3 scores, mean aggregation reduces the number of ties compared to majority voting. Similarly, our LW self-consistency methods aggregate $m$ output lists of length $D$ by minimizing Kendall-Tau distance, as described further in Sec. 3.1.2.
@@ 3.1.1 One-by-one Pointwise Methods
The first pane of Figure 1 shows one-by-one PW methods Sachan et al. (2022); Zhuang et al. (2024); Thomas et al. (2024); Upadhyay et al. (2024); Törnberg (2024); Ma et al. (2024), where each passage is scored against the query in a separate LLM call. While evaluating one candidate at a time avoids inducing candidate position biases, it also prevents the LLM from seeing potentially helpful context in other passages. Further, each passage score requires a separate LLM call, which can lead to a very large number of calls. Finally, when self-consistency ensembling is used, one-by-one methods are typically limited to repeating identical LLM calls, relying only on LLM stochasticity to generate different responses. In contrast, methods which include more than one candidate can naturally diversify prompts across an ensemble by varying co-candidate selections and permutations.
@@ 3.1.2 Listwise Ranking Methods
LW ranking methods Ma et al. (2023); Tang et al. (2024); Hou et al. (2024); Gangi Reddy et al. (2024) are shown in the second pane of Figure 1.
@@
Standard LW ranking instructs an LLM to rerank an input passage list in order of relevance with respect to $q$, with our LW prompt shown in Appendix 15. While the LLM sees all available passage context during a single inference, no absolute relevance judgments are produced.
@@
LW ranking with self-consistency Tang et al. (2024); Hou et al. (2024) involves $m$ reranking calls followed by a rank aggregation of $m$ output lists, which can be done by minimizing the Kendall-Tau distance with the $m$ lists. Multiple exact and approximate aggregation techniques exist, with our experiments using the exact Kemeny rank aggregation linear program (LP) of Tang et al. [^2] https://github.com/castorini/perm-sc We test two LW self-consistency variants:
@@ 3.2 Batched PW Methods
The right half of Figure 1 shows batched PW methods in which multiple passages are jointly scored in each LLM call. This allows an LLM to see context from multiple passages at the same time and reduces the total number of LLM calls required compared to one-by-one methods. We also study the effects of diversifying passage subsets and permutations across self-consistency ensemble prompts through several batching strategies, including all-in-one batching (c.f. Sec. 3.2.1) and sub-batching (c.f. Sec. 3.2.2). Note that autoregressively generating a sequence of scores, as opposed to only one score at a time as in one-by-one methods, could diversify scores between ensemble components further.
@@ 3.2.1 All-in-one
All-in-one PW methods prompt the LLM to score all passages in a single batch, [^3] The LLM context window must be large enough to fit all passages, otherwise sub-batching (c.f. Sec 3.2.2) is required. maximizing the context available to the model but also presenting it with the most complex task. We test two passage ordering strategies:
@@
    Shuffled: The passage list is fully shuffled before each self-consistency LLM call, giving $m$ calls with $m$ random passage permutations.
Figure 2: Effect of increasing self-consistency calls/passage ($m$) on PW relevance assessment quality (Legal Search, Shallow). For all LLMs at $m=1$ (no self-consistency), one-by-one PW is competitive, but by $m=15$ it underperforms batched PW by 5-10% AUC-PR. This shows that batching amplifies the benefits of self-consistency, likely because it allows passages to be seen jointly and creates more diverse self-consistency ensembles.
Figure 3: Effect of increasing $m$ on AUC-PR for Deep Legal Search (90 Total Psgs), showing that overly large batch sizes can harmful. The sub-batched methods (30 psgs/batch) perform very well, with the shuffling variants (STB and BTS) doing best. The large batch (90 psgs/batch) all-in-one methods perform poorly for this range of $m$, addressed further in RQ2.
@@ 3.2.2 Sub-batched
Figure 4: Median per-query runtimes for Legal Search, Shallow (30 total passages), with error bars showing the IQR. The 1st and 3rd plots show 6-17x speedups from batching. If the number of parallel LLMs available is >m>m, as in the 2nd plot, smaller batches may allow more parallelization. However, if only a few LLMs are available, processing multiple passages per call is typically much faster, especially with self-consistency ensembling.
Legal Search DL-19 Covid
PW/
LW
Psgs/
Batch
Batches/
Query
4o Sonnet Nova 4o Sonnet Nova 4o Sonnet Nova
Shallow (30)
PW 1 30 469 521 470 272 312 266 522 596 522
PW 10 3 2,956 3,212 3,032 1,074 1,208 1,082 3,502 3,968 3,569
PW 30 1 8,455 9,147 8,696 2,853 3,197 2,903 10,125 11,468 10,351
LW 30 1 8,443 9,133 8,685 2,841 3,183 2,892 10,113 11,454 10,340
Deep (90)
PW 1 90 449 500 449 271 310 265 525 599 526
PW 30 3 7,877 8,527 8,106 2,825 3,161 2,873 10,199 11,554 10,443
PW 90 1 23,179 25,063 23,892 8,094 9,045 8,267 30,268 34,286 31,136
LW 90 1 23,167 25,049 23,881 8,082 9,031 8,256 30,256 34,272 31,125
Table 1: Mean input token counts per LLM call.
@@
    Initial Order: To create $B$ batches, the initial list is partitioned into $B$ non-overlapping intervals while maintaining its order, e.g., $p_{1},\cdots,p_{30}] \rightarrow [$$p_{1}\cdots,p_{10}], [$$p_{11}\cdots,p_{20}], [$$p_{21},\cdots,p_{30}]$, where $p_{i}$ is the $i$’th passage in the initial list. Self-consistency calls for each sub-batch are identical and repeat the the same biases, giving this sub-batching strategy the least ensemble diversity.
@@
    Shuffled-then-Batched (STB): The passage list is fully shuffled and then split into $B$ batches before each LLM call, creating the most diversity in passage subsets and permutations across self-consistency ensembles.
@@
    Batched-then-Shuffled (BTS): The initial list is first divided into $B$ intervals, as in Initial Order Sub-batching above, and then each batch $b$ is fully shuffled before each LLM call. While BTS attempts to mitigate position bias through shuffling, the passage mixture for a given batch remains constant across self-consistency calls, making its LLM sampling less diverse than STB.